OpenStudy (anonymous):

How to express the congruence m^e ≡ 1 (mod n) as an expression of m?

OpenStudy (anonymous):

Has your class done logarithms yet?

OpenStudy (anonymous):

discrete logarithms? yes

OpenStudy (anonymous):

I am sorry but the last time I did this was a while back the form they want this in is m=something else correct?

OpenStudy (anonymous):

yes , modulo n

OpenStudy (anonymous):

ahhh sorry thought I had more time. I actually am at University right now and my class is about to start so very sorry if you cannot get someone else to help you out try a site called wolfram alpha It is a site capable of doing complex math and with a little playing around you may be able to get it to do what you need to here.

OpenStudy (shubhamsrg):

if m^e = 1 mod(n) then m = 1^(1/e) mod n or m= 1 modn = denotes congruence relation..

OpenStudy (anonymous):

aha, thanks anyway

OpenStudy (anonymous):

thanks. yes, but 1/e would not necessarily be an integer. the original problem is to find an integer m with given positive integers e and c such that m=e^c

OpenStudy (anonymous):

if there is such an integer, of course

OpenStudy (shubhamsrg):

8 = 2^3 ??

OpenStudy (anonymous):

yes, for example. then you give c = 8, e = 3, and it would yield 2. I need to find a test to find this number, if this exists

OpenStudy (shubhamsrg):

ohh i see..

OpenStudy (shubhamsrg):

m=e^c whats your variable here,,i mean if e and c both are integers..them m surely is an integer..

OpenStudy (anonymous):

oh. sorry, i messed up: it should be c = m^e, c and e is given, and m is to find

OpenStudy (shubhamsrg):

@mukushla might aid ,,

OpenStudy (anonymous):

Do you still need help @jsaetrum ?

OpenStudy (anonymous):

hi, yes, thanks :)